# Sample-efficient and Scalable Exploration in Continuous-Time RL

### v1

@@ C.1 GP experiments
For planning, we use the iCEM optimizer (Pinneri et al., 2021) even though it is a discrete-time algorithm. We emulate the continuous-time setting by using a fine time discretization (see measurement/control frequency $\nu$ in Table˜1) and using the equidistant MSS. The hyperparameters for the planning are given in Table˜2.
@@ C.3 Auto-tuning experiments
Here, $\bar{{\bm{\pi}}}_{n}$ denotes a slowly updated target policy obtained via Polyak averaging of ${\bm{\pi}}_{n}$. The objective promotes larger $\lambda$ values when the current policy explores less than the target policy.
@@ C.4 BNN experiments
We adopt the same hyperparameters as Sukhija et al. (2025a) for optimizing the loss via stochastic gradient descent in Equation˜8 and for configuring the UTD. We also periodically perform soft resets for the policy for training stability (D’Oro et al., 2023). The hyperparameters for the statistical model and SAC are given in Table˜3. For the ensemble-based experiments, we use several environments from the Gym and DMC benchmark suites (Brockman et al., 2016; Tunyasuvunakool et al., 2020). We adapt them to the continuous-time setting by approximating the derivatives using a finite difference filter. The measurement/control frequency is given by the duration of an environment step dt.
@@
Environments    Action Repeat        Policy / Critic Architecture        Model Architecture        Learning Rate       Batch Size Gym – Pendulum / MountainCar 1 (256,256) $5\times$(256,256) $3\times 10^{-4}$ 256 Gym – Reacher / DMC – Quadruped 2 (256,256) $5\times$(512,512) $3\times 10^{-4}$ 256 Gym – other environmentsa 2 (256,256) $5\times$(256,256) $3\times 10^{-4}$ 256 DMC – all except Quadruped 2 (256,256) $5\times$(256,256) $3\times 10^{-4}$ 256 a HalfCheetah, Hopper, Pusher.
@@ C.5 Downstream tasks
C.6 Time-adaptive experiments
@@
For the time-adaptive experiments, we evaluate COMBRL on two continuous-time environments. We first reuse the Pendulum-GP environment described in the GP experiments (see Section˜C.1), with an added interaction cost $C$.
@@
We give the relevant parameters for the time-adaptive experiments in Table˜5. In Figure˜5, we show the final performance at convergence of the evaluated algorithms. For completeness, we also offer the complete learning curves in Figure˜6.

### v2

@@ C.1 GP experiments
For planning, we use the iCEM optimizer (Pinneri et al., 2021) even though it is a discrete-time algorithm. The method iteratively samples action sequences, evaluates them in the dynamics model, and refits a Gaussian distribution to the top-performing (“elite”) trajectories. This elite set determines the parameters for the next iteration.
In contrast to vanilla CEM introduced by Rubinstein (1999), iCEM improves efficiency by sampling temporally correlated action sequences and reusing elite trajectories across iterations and MPC steps. We emulate the continuous-time setting by using a fine time discretization (see measurement/control frequency $\nu$ in Table˜1) and using the equidistant MSS. The hyperparameters for the planning are given in Table˜2.
@@ C.3 Auto-tuning experiments
Here, $\bar{{\bm{\pi}}}_{n}$ denotes a slowly updated target policy obtained via Polyak averaging of ${\bm{\pi}}_{n}$. The objective promotes larger $\lambda$ values when the current policy explores less than the target policy. Figure˜6 shows that the auto-tuned intrinsic reward weight $\lambda_{n}$ typically decreases over training for several of the tasks shown in Figure˜2, indicating that exploration is emphasized early on and gradually reduced as the model and policy stabilize.
Figure 6: Evolution of the auto-tuned intrinsic reward weight $\lambda_{n}$. We plot $\lambda_{n}$ as a function of environment steps for COMBRL with the auto-tuning objective in Equation˜8 across several Gym/DMC tasks, as seen in Figure˜2. Overall, $\lambda_{n}$ tends to decrease over time, indicating stronger exploration early in training and reduced weighting on the intrinsic reward as learning progresses.
@@ C.4 BNN experiments
The policy is optimized using soft actor-critic (SAC, Haarnoja et al., 2018), an off-policy actor-critic method that learns a stochastic policy by maximizing expected return together with an entropy bonus to encourage exploration and stabilize training. We adopt the same hyperparameters as Sukhija et al. (2025a) for optimizing the loss via stochastic gradient descent in Equation˜8 and for configuring the UTD. We also periodically perform soft resets for the policy for training stability (D’Oro et al., 2023). The hyperparameters for the statistical model and SAC are given in Table˜3. For the ensemble-based experiments, we use several environments from the Gym and DMC benchmark suites (Brockman et al., 2016; Tunyasuvunakool et al., 2020). We adapt them to the continuous-time setting by approximating the derivatives using a finite difference filter. The measurement/control frequency is given by the duration of an environment step dt.
@@
Environments    Action Repeat        Policy / Critic Architecture        Model Architecture        Learning Rate       Batch Size Gym – Pendulum / MountainCar 1 (256,256) $5\times$(256,256) $3\times 10^{-4}$ 256 Gym – Reacher 2 (256,256) $5\times$(512,512) $3\times 10^{-4}$ 256 Gym – other environmentsa 2 (256,256) $5\times$(256,256) $3\times 10^{-4}$ 256 DMC – Quadruped 2 (256,256) $5\times$(512,512) $3\times 10^{-4}$ 256 DMC – Humanoid 2 (512,512) $5\times$(512,512) $3\times 10^{-4}$ 256 DMC – other environmentsb 2 (256,256) $5\times$(256,256) $3\times 10^{-4}$ 256 a HalfCheetah, Hopper, Pusher. b Hopper, CartPole.
@@ C.5 Downstream tasks
C.6 Epistemic uncertainty decrease
We visualize how COMBRL reduces epistemic uncertainty over time in both the unsupervised and reward-driven regimes in Figure˜7.
In the unsupervised setting, we run COMBRL with a purely intrinsic objective (i.e., $\lambda_{n}\to\infty$) on the Pendulum environment and monitor epistemic uncertainty over the entire reachable state-action space. Concretely, at the end of each episode $n$ we sample 100 random state-action pairs ${\bm{z}}=({\bm{x}},{\bm{u}})$ from ${\mathcal{X}}\times{\mathcal{U}}$ and evaluate the model uncertainty $\|{\bm{\sigma}}_{n}({\bm{z}})\|$ at these points. This shows that unsupervised exploration alone is sufficient to drive down epistemic uncertainty globally over ${\mathcal{X}}\times{\mathcal{U}}$, rather than only along the visited trajectories.
In the reward-driven setting with auto-tuned $\lambda_{n}$, we consider the CartPole environment and use the method from Appendix C.3 to auto-tune $\lambda_{n}$. Here we track the epistemic uncertainty encountered along the executed trajectories at each environment step. Compared to the Mean-planning and PETS baselines, COMBRL consistently visits regions with higher epistemic uncertainty, indicating that the intrinsic reward term steers the agent towards informative parts of the state-action space, while the overall level of uncertainty still decreases over time.
Figure 7: Epistemic uncertainty under COMBRL. Left: Unsupervised COMBRL ($\lambda_{n}\to\infty$) on Pendulum. At the end of each episode, we evaluate the model uncertainty $\|{\bm{\sigma}}_{n}(z)\|$ on 100 random samples $z\in{\mathcal{X}}\times{\mathcal{U}}$ and track its evolution, showing that unsupervised exploration systematically reduces epistemic uncertainty across the entire reachable state-action space. Right: COMBRL with auto-tuned intrinsic reward weight $\lambda_{n}$ on DMC-CartPole. We plot the epistemic uncertainty encountered along the executed trajectories at each environment step. Compared to Mean and PETS, COMBRL is guided towards regions of higher epistemic uncertainty while still driving uncertainty down over time, indicating that $\lambda_{n}$ effectively shapes exploration towards informative state-action regions.
C.7 Time-adaptive experiments
@@
For the time-adaptive experiments, we evaluate COMBRL on two continuous-time environments. We first reuse the Pendulum-GP environment described in the GP experiments (see Appendix C.1), with an added interaction cost $C$.
@@
We give the relevant parameters for the time-adaptive experiments in Table˜5. In Figure˜5, we show the final performance at convergence of the evaluated algorithms. For completeness, we also offer the complete learning curves in Figure˜8.